OpenStudy (anonymous):

use factoring to solve this equation x^2-8x+15=0

OpenStudy (anonymous):

x= 8 +- 2 ---------- 2

OpenStudy (anonymous):

product = 15, sum = -8..

OpenStudy (anonymous):

x^2-8x+15=0 x^2 -5x -3x + 15 = 0 x(x-5) -3(x-5) = 0 (x-3)(x-5) = 0 this can be true only if either x-3=0 or x-5=0 if x-3 = 0 => x=3 if x-5 = 0 => x=5 solutions set is x = 3,5